media%2F4d1%2F4d14ee5f-070a-4df4-9b11-3d
Answer #1

First need to calculate the ΔH and is a product of heat capacity (Cp) and temperature change (ΔT)

ΔH=418x1.48=618.64 J/K

Now estimate the number of moles of ATP mixed with ATPase

You have taken One hundred milliliters of 0.200 M ATP, so moles= (100mlx1L/1000ml)x0.2 mol/L=0.02 mol

What is ΔH for the hydrolysis of ATP to ADP and phosphate= ΔH/mol=618.64 J/K/0.02 mol=30,932 Joules/mol or 30.932 kJ/mol